# Direct Preference Optimization Needs

### v3

@@ 4 Experiments
$\alpha$-DPO 58.7 51.1 30.8 36.3 35.7 73.4 66.1 48.6 59.3 60.8
In this section, we present the main results of our experiments, highlighting the superior performance of $\alpha$-DPO over existing methods on various benchmarks and ablation studies to analyze the impact of different components of $\alpha$-DPO.
@@ 4.1 Experiments Setup
Table 2: Ablation studies under Llama3-Instruct v0.2 and Mistral-Instruct settings. We ablate each key design of $\alpha$-DPO and explore variants of the implicit reference model $\hat{\pi}_{\text{ref}}$.
@@
Baselines. We compare $\alpha$-DPO with several state-of-the-art preference optimization methods: DPO (Rafailov et al., 2023), SimPO (Meng et al., 2024), IPO (Azar et al., 2023), CPO (Xu et al., 2024), KTO (Ethayarajh et al., 2024), ORPO (Hong et al., 2024), and R-DPO (Park et al., 2024). We also include the SFT model as a baseline. We thoroughly tune the hyperparameters for each baseline and report the best performance. Further details can be found in Appendix B.1.
@@ 4.2 Main Results
$\alpha$-DPO consistently outperforms existing preference optimization methods. As shown in Table 1, while all preference optimization algorithms improve over the SFT baseline, $\alpha$-DPO achieves superior performance compared to existing methods specifically on the AlpacaEval 2 LC metric. These significant improvements highlight the robustness and effectiveness of $\alpha$-DPO. Specifically, $\alpha$-DPO outperforms the best baseline by an average of 3 percentage points in AlpacaEval 2 LC win rate. Furthermore, on benchmarks such as Arena-Hard, $\alpha$-DPO achieves state-of-the-art or second-best results, demonstrating its competitiveness across different evaluation settings.
@@
The importance of the design on the implicit reference model. As the core contribution of this work is to propose a novel reference model $\hat{\pi}_{\text{ref}}(y|x)=U(y|x)\left({\pi_{\theta}(y|x)}/{\pi_{\text{ref}}(y|x)}\right)^{\alpha}$, we also evaluate other variants of the reference model. Specifically, we compare $\alpha$-DPO with three variants: (1) $\hat{\pi}_{\text{ref}}(y|x)=U(y|x)\left({\pi_{\theta}(y|x)}\right)^{\alpha}$, (2) $\hat{\pi}_{\text{ref}}(y|x)=U(y|x)\left({\pi_{\text{ref}}(y|x)}\right)^{\alpha}$, and (3) $\hat{\pi}_{\text{ref}}(y|x)=U(y|x)\left(1/{\pi_{\text{ref}}(y|x)}\right)^{\alpha}$. As shown in Table 2, most of the variants perform better than SimPO ($\hat{\pi}_{\text{ref}}(y|x)=U(y|x)$), which demonstrates the importance of adaptive margin between pairs. Besides, our proposed reference model consistently outperforms other variants, indicating the effectiveness of the proposed design.
All key designs in $\alpha$-DPO are crucial. To further analyze the impact of different components in $\alpha$-DPO, we conduct ablation studies by removing key components from $\alpha$-DPO. As shown in Table 2, removing normalization, stop gradient, or setting $\gamma=0$ all lead to significant performance drops, highlighting the importance of these components in $\alpha$-DPO.
4.3 KL divergence control in $\alpha$-DPO
Outstanding Performance and Lower KL. As noted in Rafailov et al. (2023); Zeng et al. (2024), it is crucial to consider both performance and KL divergence when comparing algorithms. A slightly higher win rate accompanied by a significantly higher KL divergence is often not desirable. In line with the design principles of TDPO, we implemented SimPO and $\alpha$-DPO. Figure 2(a) 2(b) presents the KL divergence curves. The results indicate that as $\alpha$ increases, the KL divergence of $\alpha$-DPO remains stable or even decreases slightly when compared to SimPO. This demonstrates that $\alpha$-DPO not only achieves superior performance but also maintains a lower KL divergence, indicating a better balance between alignment and control of KL divergence during the training process.
Mitigating Over-Optimization. Over-optimization, as described by Gao et al. (2023) and Rafailov et al. (2024), refers to a phenomenon where model performance exhibits a hump-shaped pattern across different targets: beyond an optimal point, further increasing the KL budget results in diminishing returns. To investigate this, we evaluate SimPO and $\alpha$-DPO at four intermediate checkpoints, corresponding to different KL budgets. As illustrated in Figure 2(c), it is intriguing that while the performance of our approach does decrease with increasing KL budget, the decline is relatively modest. This indicates that our method effectively mitigates the issue of over-optimization.
@@ 4.3 KL divergence control in $\alpha$-DPO
Figure 3: Impact of $\alpha$ on (a) LC and SC win rate, (b) reward difference distribution, and (c) log-likelihood distribution of chosen responses in $\alpha$-DPO.
4.4 The Impact of $\alpha$ in $\alpha$-DPO
@@
Effect of $\alpha$ on Performance. We investigated how the parameter $\alpha$ in $\alpha$-DPO impacts the win rate on AlpacaEval 2 and Arena-Hard. The results, as shown in Figure 3 (a), indicate that the style-control win rate on Arena-Hard initially increases and then decreases with increasing $\alpha$. In contrast, the length-control win rate on AlpacaEval 2 exhibits a consistently increasing trend. This suggests that the optimal value of $\alpha$ varies depending on the evaluation benchmarks. Further experiments refer to Appendix B.3.
@@
Impact of $\alpha$ on the reward distribution. We visualize the distribution of the learned reward margin $r(x,y_{w})-r(x,y_{l})$ and the log likelihood of the chosen response $\log\pi_{\theta}(y_{w}|x)$ under different $\alpha$ values in Figure 3 (b,c). Decreasing $\alpha$ results in a flatter reward margin, while the log likelihood distribution remains relatively unchanged. Conversely, in SimPO (cf. Figure 5), increasing $\gamma$ yields a flatter reward margin distribution but at the cost of also flattening the log likelihood distribution, which undesirably lowers the log likelihood of positive samples. This indicates that $\alpha$-DPO can better balance the relationship between the reward margin and log likelihood.
5 Related Work
Reinforcement learning from human feedback. RLHF is a technique that aligns large language models with human preferences and values (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022; Azar et al., 2023). Traditional RLHF can be divided into three stages: supervised fine-tuning (Zhou et al., 2023; Taori et al., 2023; Geng et al., 2023; Conover et al., 2023; Köpf et al., 2023; Ding et al., 2023), reward modeling (Gao et al., 2023; Luo et al., 2023; Chen et al., 2024b; Lightman et al., 2023; Havrilla et al., 2024; Lambert et al., 2024), and policy optimization (Schulman et al., 2017; Anthony et al., 2017). In the third stage, Proximal Policy Optimization (PPO) is a widely used algorithm. Additionally, Xiong et al. (2023) proposed efficient algorithms for the reverse-KL regularized contextual bandit framework in RLHF. Ye et al. (2024) introduced provably efficient algorithms for KL-regularized Nash-Learning from Human Feedback (NLHF). Furthermore, Ji et al. (2024b) developed an active-query-based PPO algorithm with specific regret bounds and query complexity.
Offline direct preference optimization. Several alternative preference optimization objectives have been proposed in addition to DPO (Rafailov et al., 2023). IPO (Azar et al., 2023) addresses the overfitting issues associated with DPO. ORPO (Hong et al., 2024) and SimPO (Meng et al., 2024) aim to eliminate the dependence on a reference model. R-DPO (Park et al., 2024) focuses on mitigating exploitation based on sequence length. KTO (Ethayarajh et al., 2024) deals with preference optimization when data are not pairwise. CPO (Xu et al., 2024) and $\beta$-DPO(Wu et al., 2024) emphasize the quality of preference data. Another line of research explores comparisons among more than two instances (Dong et al., 2023; Liu et al., 2024; Song et al., 2024; Yuan et al., 2023).
Online direct preference optimization. Offline direct preference optimization methods are simple but rely on preference data collected offline. RLHF methods interact online with the language model being aligned but require policy gradients. In contrast, online direct preference optimization methods combine the advantages of both approaches. Yuan et al. (2024) proposed a “self-rewarding” approach in which the policy being aligned provides online feedback to itself. Alternatively, OAIF (Guo et al., 2024) is a novel online preference optimization method that can leverage feedback from any LLM, including those stronger than the LLM being aligned. Swamy et al. (2024) also concurrently investigate the importance of online preference but still rely on reward models (RMs). SELMA (Zhang et al., 2024) improves exploration efficiency by selectively favoring responses with high potential rewards rather than indiscriminately sampling unseen responses.
@@ 6 Discussion
Conclusion. We proposed $\alpha$-DPO, an adaptive preference optimization method that improves LLM alignment by introducing a dynamic reward margin based on instance-specific differences. $\alpha$-DPO addresses limitations in previous methods like DPO and SimPO by balancing alignment and diversity through KL divergence control. Our theoretical guarantees and empirical results show that $\alpha$-DPO consistently outperforms baselines on benchmarks like AlpacaEval 2 and Arena-Hard, with significant improvements in win rates, establishing it as a robust solution for LLM fine-tuning.
@@
Limitations and Future Work. While $\alpha$-DPO enhances performance, it introduces an additional hyperparameter, $\alpha$, requiring manual tuning. Future work could focus on developing an adaptive approach to automatically adjust this parameter. Additionally, although we show $\alpha$-DPO’s theoretical equivalence to online methods, it remains an offline approach. Extending it to online learning would allow real-time adaptation, broadening its application in interactive environments. Lastly, we observed that different benchmarks, such as AlpacaEval 2 and Arena-Hard, require distinct parameter settings for optimal performance. Investigating a more generalized approach that adapts effectively across multiple benchmarks would further improve the model’s versatility.
Appendix A Appendix
A.1 Proof of Theorem 2.1
@@ Appendix A Appendix
A.2 Proof of Lemma 3.2
See 3.2
Proof.
We aim to establish an inequality between the losses $L_{1}=\mathcal{L}_{\text{SimPO}}^{\text{online}}(\pi_{\theta},\pi_{\text{ref}})$ and $L_{2}=\mathcal{L}_{\text{$\alpha$-DPO}}(\pi_{\theta},\pi_{\text{ref}})$.
Recall that the online SimPO loss $L_{1}$ is given by:
$L_{1}=-\mathbb{E}_{\pi_{\text{ref}}}\left[\log\sigma\left(\beta\log\pi_{\theta}(y_{w}|x)-\beta\log\pi_{\theta}(y_{l}|x)-\gamma\right)\cdot w(y_{w},y_{l}|x)\right],$
where the importance weight $w(y_{w},y_{l}|x)$ is:
$w(y_{w},y_{l}|x)=\frac{\pi_{\theta,\text{old}}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}\cdot\frac{\pi_{\text{ref}}(y_{l}|x)}{\pi_{\theta,\text{old}}(y_{l}|x)}.$
Under the assumption that the old policy $\pi_{\theta,\text{old}}$ is related to the current policy $\pi_{\theta}$ via:
$\pi_{\theta,\text{old}}(y|x)=C(y|x)\left(\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}\right)^{\alpha}\pi_{\text{ref}}(y|x),$
with $C(y|x)$ being a normalization constant, we can express the importance weight as:
$w(y_{w},y_{l}|x)$ $=\frac{C(y_{w}|x)\left(\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}\right)^{\alpha}\pi_{\text{ref}}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}\cdot\frac{\pi_{\text{ref}}(y_{l}|x)}{C(y_{l}|x)\left(\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)^{\alpha}\pi_{\text{ref}}(y_{l}|x)}$
$=\frac{C(y_{w}|x)}{C(y_{l}|x)}\left(\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}\cdot\frac{\pi_{\text{ref}}(y_{l}|x)}{\pi_{\theta}(y_{l}|x)}\right)^{\alpha}.$
Taking the logarithm, we have:
$\log w(y_{w},y_{l}|x)=\log\frac{C(y_{w}|x)}{C(y_{l}|x)}+\alpha B,$
where
$B=\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}.$
For sufficiently small $\alpha$, we can approximate the importance weight using the first-order Taylor expansion:
$w(y_{w},y_{l}|x)=\exp(\log w(y_{w},y_{l}|x))\approx\exp(\alpha B)=1+\alpha B.$
Thus, the loss $L_{1}$ can be approximated as:
$L_{1}$ $\approx-\mathbb{E}_{\pi_{\text{ref}}}\left[\log\sigma(A)\left(1+\alpha B\right)\right]$
$=-\mathbb{E}_{\pi_{\text{ref}}}\left[\log\sigma(A)\right]-\alpha\mathbb{E}_{\pi_{\text{ref}}}\left[B\log\sigma(A)\right],$
where
$A=\beta\log\pi_{\theta}(y_{w}|x)-\beta\log\pi_{\theta}(y_{l}|x)-\gamma.$
Meanwhile, the $\alpha$-DPO loss $L_{2}$ is given by:
$L_{2}=-\mathbb{E}_{\pi_{\text{ref}}}\left[\log\sigma\left(A-\alpha B\right)\right].$
Expanding $\log\sigma(A-\alpha B)$ around $\alpha=0$ using the first-order Taylor expansion, we obtain:
$\log\sigma(A-\alpha B)\approx\log\sigma(A)-\alpha B\left(1-\sigma(A)\right),$
since $\frac{d}{dA}\log\sigma(A)=1-\sigma(A)$.
Therefore, $L_{2}$ becomes:
$L_{2}\approx-\mathbb{E}_{\pi_{\text{ref}}}\left[\log\sigma(A)-\alpha B\left(1-\sigma(A)\right)\right]=-\mathbb{E}_{\pi_{\text{ref}}}\left[\log\sigma(A)\right]+\alpha\mathbb{E}_{\pi_{\text{ref}}}\left[B\left(1-\sigma(A)\right)\right].$
Subtracting $L_{2}$ from $L_{1}$, we have:
$L_{1}-L_{2}$ $\approx\left(-\mathbb{E}_{\pi_{\text{ref}}}\left[\log\sigma(A)\right]-\alpha\mathbb{E}_{\pi_{\text{ref}}}\left[B\log\sigma(A)\right]\right)-\left(-\mathbb{E}_{\pi_{\text{ref}}}\left[\log\sigma(A)\right]+\alpha\mathbb{E}_{\pi_{\text{ref}}}\left[B\left(1-\sigma(A)\right)\right]\right)$
$=-\alpha\mathbb{E}_{\pi_{\text{ref}}}\left[B\log\sigma(A)\right]-\alpha\mathbb{E}_{\pi_{\text{ref}}}\left[B\left(1-\sigma(A)\right)\right]$
$=-\alpha\mathbb{E}_{\pi_{\text{ref}}}\left[B\left(\log\sigma(A)+1-\sigma(A)\right)\right].$
Rewriting the above, we get:
$L_{2}\leq L_{1}+\alpha\mathbb{E}_{\pi_{\text{ref}}}\left[B\left(\log\sigma(A)-\sigma(A)+1\right)\right].$
Thus, we have established that:
$\mathcal{L}_{\text{$\alpha$-DPO}}(\pi_{\theta},\pi_{\text{ref}})\leq\mathcal{L}_{\text{SimPO}}^{\text{online}}(\pi_{\theta},\pi_{\text{ref}})+\mathbb{E}_{\pi_{\text{ref}}}\left[\alpha B\left(\log\sigma(A)-\sigma(A)+1\right)\right],$
which completes the proof. ∎
A.3 Proof of Lemma 3.3
See 3.3
@@
where $M(x,y_{w},y_{l})$ is the margin term defined in the $\alpha$-DPO objective. Thus, we have shown that:
@@
Appendix B Experiments
B.1 Implementation Details

### v4

@@ 4 Experiments
AlphaDPO 58.7 51.1 30.8 36.3 35.7 73.4 66.1 48.6 59.3 60.8
In this section, we present the main results of our experiments, highlighting the superior performance of AlphaDPO over existing methods on various benchmarks and ablation studies to analyze the impact of different components of AlphaDPO.
@@ 4.1 Experiments Setup
Table 2: Ablation studies under Llama3-Instruct v0.2 and Mistral-Instruct settings. We ablate each key design of AlphaDPO and explore variants of the implicit reference model $\hat{\pi}_{\text{ref}}$.
@@
Baselines. We compare AlphaDPO with several state-of-the-art preference optimization methods: DPO (Rafailov et al., 2023), SimPO (Meng et al., 2024), IPO (Azar et al., 2024), CPO (Xu et al., 2024), KTO (Ethayarajh et al., 2024), ORPO (Hong et al., 2024), and R-DPO (Park et al., 2024). We also include the SFT model as a baseline. We thoroughly tune the hyperparameters for each baseline and report the best performance. Further details can be found in Appendix D.1.
@@ 4.2 Main Results
AlphaDPO consistently outperforms existing preference optimization methods. As shown in Table 1, while all preference optimization algorithms improve over the SFT baseline, AlphaDPO achieves superior performance compared to existing methods specifically on the AlpacaEval 2 LC metric. These significant improvements highlight the robustness and effectiveness of AlphaDPO. Specifically, AlphaDPO outperforms the best baseline by an average of 3 percentage points in AlpacaEval 2 LC win rate. Furthermore, on benchmarks such as Arena-Hard, AlphaDPO achieves state-of-the-art or second-best results, demonstrating its competitiveness across different evaluation settings.
@@
The importance of the design on the implicit reference model. As the core contribution of this work is to propose a novel reference model $\hat{\pi}_{\text{ref}}(y|x)=U(y|x)\left({\pi_{\theta}(y|x)}/{\pi_{\text{ref}}(y|x)}\right)^{\alpha}$, we also evaluate other variants of the reference model. Specifically, we compare AlphaDPO with three variants: (1) $\hat{\pi}_{\text{ref}}(y|x)=U(y|x)\left({\pi_{\theta}(y|x)}\right)^{\alpha}$, (2) $\hat{\pi}_{\text{ref}}(y|x)=U(y|x)\left({\pi_{\text{ref}}(y|x)}\right)^{\alpha}$, and (3) $\hat{\pi}_{\text{ref}}(y|x)=U(y|x)\left(1/{\pi_{\text{ref}}(y|x)}\right)^{\alpha}$. As shown in Table 2, most of the variants perform better than SimPO ($\hat{\pi}_{\text{ref}}(y|x)=U(y|x)$), which demonstrates the importance of adaptive margin between pairs. Besides, our proposed reference model consistently outperforms other variants, indicating the effectiveness of the proposed design.
All key designs in AlphaDPO are crucial. To further analyze the impact of different components in AlphaDPO, we conduct ablation studies by removing key components from AlphaDPO. As shown in Table 2, removing normalization, stop gradient, or setting $\gamma=0$ all lead to significant performance drops, highlighting the importance of these components in AlphaDPO.
4.3 KL divergence control in AlphaDPO
Outstanding Performance and Lower KL. As noted in Rafailov et al. (2023); Zeng et al. (2024), it is crucial to consider both performance and KL divergence when comparing algorithms. A slightly higher win rate accompanied by a significantly higher KL divergence is often not desirable. In line with the design principles of TDPO, we implemented SimPO and AlphaDPO. Figure 2(a) 2(b) presents the KL divergence curves. The results indicate that as $\alpha$ increases, the KL divergence of AlphaDPO remains stable or even decreases slightly when compared to SimPO. This demonstrates that AlphaDPO not only achieves superior performance but also maintains a lower KL divergence, indicating a better balance between alignment and control of KL divergence during the training process.
@@ 4.3 KL divergence control in AlphaDPO
Figure 3: Impact of $\alpha$ on (a) LC and SC win rate, (b) reward difference distribution, and (c) log-likelihood distribution of chosen responses in AlphaDPO.
Mitigating Over-Optimization. Over-optimization, as described by Gao et al. (2023) and Rafailov et al. (2024), refers to a phenomenon where model performance exhibits a hump-shaped pattern across different targets: beyond an optimal point, further increasing the KL budget results in diminishing returns. To investigate this, we evaluate SimPO and AlphaDPO at four intermediate checkpoints, corresponding to different KL budgets. As illustrated in Figure 2(c), it is intriguing that while the performance of our approach does decrease with increasing KL budget, the decline is relatively modest. This indicates that our method effectively mitigates the issue of over-optimization.
4.4 The Impact of $\alpha$ in AlphaDPO
@@
Effect of $\alpha$ on Performance. We investigated how the parameter $\alpha$ in AlphaDPO impacts the win rate on AlpacaEval 2 and Arena-Hard. The results, as shown in Figure 3 (a), indicate that the style-control win rate on Arena-Hard initially increases and then decreases with increasing $\alpha$. In contrast, the length-control win rate on AlpacaEval 2 exhibits a consistently increasing trend. This suggests that the optimal value of $\alpha$ varies depending on the evaluation benchmarks. Further experiments refer to Appendix D.3.
@@
Impact of $\alpha$ on the reward distribution. We visualize the distribution of the learned reward margin $r(x,y_{w})-r(x,y_{l})$ and the log likelihood of the chosen response $\log\pi_{\theta}(y_{w}|x)$ under different $\alpha$ values in Figure 3 (b,c). Decreasing $\alpha$ results in a flatter reward margin, while the log likelihood distribution remains relatively unchanged. Conversely, in SimPO (cf. Figure 5), increasing $\gamma$ yields a flatter reward margin distribution but at the cost of also flattening the log likelihood distribution, which undesirably lowers the log likelihood of positive samples. This indicates that AlphaDPO can better balance the relationship between the reward margin and log likelihood.
@@ 5 Discussion
Conclusion. We proposed AlphaDPO, an adaptive preference optimization method that improves LLM alignment by introducing a dynamic reward margin based on instance-specific differences. AlphaDPO addresses limitations in previous methods like DPO and SimPO by balancing alignment and diversity through KL divergence control. Our theoretical guarantees and empirical results show that AlphaDPO consistently outperforms baselines on benchmarks like AlpacaEval 2 and Arena-Hard, with significant improvements in win rates, establishing it as a robust solution for LLM fine-tuning.
@@
Limitations and Future Work. While AlphaDPO enhances performance, it introduces an additional hyperparameter, $\alpha$, requiring manual tuning. Future work could focus on developing an adaptive approach to automatically adjust this parameter. Additionally, although we show AlphaDPO’s theoretical equivalence to online methods, it remains an offline approach. Extending it to online learning would allow real-time adaptation, broadening its application in interactive environments. Lastly, we observed that different benchmarks, such as AlpacaEval 2 and Arena-Hard, require distinct parameter settings for optimal performance. Investigating a more generalized approach that adapts effectively across multiple benchmarks would further improve the model’s versatility.
Impact Statement
This work introduces AlphaDPO, an adaptive preference optimization framework that improves Large Language Model (LLM) alignment with human preferences. By establishing a more principled and effective mechanism for preference optimization, AlphaDPO contributes to developing more robustly aligned LLMs, a crucial step towards safer and more beneficial AI systems.
Acknowledgment
This research is supported by the National Science and Technology Major Project (2023ZD0121102), and the National Natural Science Foundation of China (62302321, U24B20180). This research was also supported by the advanced computing resources provided by the Supercomputing Center of the USTC.
Appendix A Related Work
Reinforcement learning from human feedback. RLHF is a technique that aligns large language models with human preferences and values (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022; Azar et al., 2024). Traditional RLHF can be divided into three stages: supervised fine-tuning (Zhou et al., 2023; Taori et al., 2023; Geng et al., 2023; Conover et al., 2023; Köpf et al., 2023; Ding et al., 2023), reward modeling (Gao et al., 2023; Luo et al., 2025; Chen et al., 2024b; Lightman et al., 2024; Havrilla et al., 2024; Lambert et al., 2025), and policy optimization (Schulman et al., 2017; Anthony et al., 2017). In the third stage, Proximal Policy Optimization (PPO) is a widely used algorithm. Additionally, Xiong et al. (2023) proposed efficient algorithms for the reverse-KL regularized contextual bandit framework in RLHF. Ye et al. (2024) introduced provably efficient algorithms for KL-regularized Nash-Learning from Human Feedback (NLHF). Furthermore, Ji et al. (2024b) developed an active-query-based PPO algorithm with specific regret bounds and query complexity.
Offline direct preference optimization. Several alternative preference optimization objectives have been proposed in addition to DPO (Rafailov et al., 2023). IPO (Azar et al., 2024) addresses the overfitting issues associated with DPO. ORPO (Hong et al., 2024) and SimPO (Meng et al., 2024) aim to eliminate the dependence on a reference model. R-DPO (Park et al., 2024) focuses on mitigating exploitation based on sequence length. KTO (Ethayarajh et al., 2024) deals with preference optimization when data are not pairwise. CPO (Xu et al., 2024) and $\beta$-DPO(Wu et al., 2024) emphasize the quality of preference data. Another line of research explores comparisons among more than two instances (Dong et al., 2023; Liu et al., 2025a; Song et al., 2024; Yuan et al., 2023).
Online direct preference optimization. Offline direct preference optimization methods are simple but rely on preference data collected offline. RLHF methods interact online with the language model being aligned but require policy gradients. In contrast, online direct preference optimization methods combine the advantages of both approaches. Yuan et al. (2024) proposed a “self-rewarding” approach in which the policy being aligned provides online feedback to itself. Alternatively, OAIF (Guo et al., 2024) is a novel online preference optimization method that can leverage feedback from any LLM, including those stronger than the LLM being aligned. Swamy et al. (2024) also concurrently investigate the importance of online preference but still rely on reward models (RMs). SELMA (Zhang et al., 2025) improves exploration efficiency by selectively favoring responses with high potential rewards rather than indiscriminately sampling unseen responses.
Appendix B Proofs
B.1 Proof of Theorem 2.1
@@ Appendix B Proofs
B.2 Proof of Lemma 3.1
See 3.1
@@
where $M(x,y_{w},y_{l})$ is the margin term defined in the AlphaDPO objective. Thus, we have shown that:
@@
Appendix C The motivation for the proposed $\hat{\pi}_{\text{ref}}(y|x)$
The motivation for the proposed reference policy $\hat{\pi}_{\text{ref}}(y|x)$ can be clarified as follows:
  • Utility Theory Perspective: The proposed $\hat{\pi}_{\text{ref}}(y|x)$ is designed with the uniform distribution $U(y|x)$ as a baseline. The term $\left(\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}\right)^{\alpha}$ dynamically adjusts the reward margin by balancing contributions from the policy and reference models. This mechanism can be interpreted through the lens of utility theory as relative attractiveness, enabling adaptive instance-specific reward modeling.
  • Gradient Perspective By introducing $\hat{\pi}_{\text{ref}}(y|x)$, the framework mitigates the label flipping issues found in DPO or SimPO. In the SimPO framework, the gradient is expressed as:
    $\nabla_{\theta}\mathcal{L}_{\mathrm{SimPO}}(\pi_{\theta})=-\beta\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[s_{\theta}\left(\frac{1}{|y_{w}|}\nabla_{\theta}\log\pi_{\theta}(y_{w}|x)-\frac{1}{|y_{l}|}\nabla_{\theta}\log\pi_{\theta}(y_{l}|x)\right)\right],$
    where $s_{\theta}=\sigma\left(\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)+\gamma\right)$.
    This formulation may amplify weights when the reward estimate is incorrect. By contrast, under AlphaDPO:
    $s_{\theta}=\sigma\left(\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)+\gamma+\alpha M(x,y_{w},y_{l})\right),$
    the additional $\alpha M(x,y_{w},y_{l})$ component increases weight when the reward estimate is accurate, ensuring a more robust reward signal.
  • Motivational Core The central goal of the proposed AlphaDPO is to address the unreliability of the reference policy, as outlined in Section 3.1. By integrating the policy model into the reference model design, the quality of the reference model is enhanced, improving fine-tuning performance. Similar concepts have been explored in recent works (Gorbatovski et al., 2025; Liu et al., 2025b).
Appendix D Experiments
D.1 Implementation Details